OpenStudy (firejay5):

Simplify. Multiply Pure Imaginary Numbers. 22.) the square root of -13 * the square root of -26.

OpenStudy (raden):

sqrt(-13) = sqrt(-1*13) = sqrt(-1) sqrt(13) = i sqrt(13) sqrt(-26) = sqrt(-1*2*13) = sqrt(-1) sqrt(2) sqrt(13) = i sqrt(2) sqrt(13) so, multiply of them : sqrt(-13) * sqrt(-26) = i sqrt(13) * i sqrt(2) sqrt(13) continue it....

OpenStudy (firejay5):

are you sure that's correct

OpenStudy (raden):

yes, still incomplete :) can u simplify it

OpenStudy (firejay5):

Let's see stay on here for awhile not long

OpenStudy (firejay5):

so is your work written like this: \[i \sqrt{13} * i \sqrt{2} * \sqrt{13}\]

OpenStudy (raden):

yes, now simplify it ....

OpenStudy (firejay5):

\[13i \sqrt{2}\]

OpenStudy (raden):

i * i = i^1 = (sqrt(-1))^2 = -1 the rest have corrected

OpenStudy (firejay5):

so what did I do wrong

OpenStudy (raden):

i^2 i mean :)

OpenStudy (raden):

cancel out i, because has be -1

OpenStudy (firejay5):

so would it be -13

OpenStudy (raden):

where's the sqrt(2) ?

OpenStudy (firejay5):

\[-13\sqrt{2}\]

OpenStudy (raden):

thats all right ^^

OpenStudy (firejay5):

so there is no i in this problem

OpenStudy (raden):

i was death :p

OpenStudy (firejay5):

It's basically because it got cancelled out right or what

OpenStudy (raden):

just understanding define of i = sqrt(-1) so, i * i = sqrt(-1) * sqrt(-1) = {sqrt(-1)}^2 = -1 as proportion, let's we ask to the best calculator : http://www.wolframalpha.com/input/?i=sqrt%28-13%29sqrt%28-26%29